# Affine Chord Sobolev Inequalities from Integral Geometry

### v1

@@ 1. Introduction
Using the second correspondence, Alonso-Gutiérrez, Bernués and González Merino [2] extended Zhang’s inequality to log-concave functions, showing that for a log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy\leq 2^{n}n!\|f\|_{1}^{n+1}|\Pi^{*}f|.$ (1.11)
If $f(0)=\|f\|_{\infty}$, equality holds if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$. For further results on Zhang’s inequality and on radial mean bodies, we refer the reader to [3, 4, 20, 21, 24, 28].
In Section 9, we extend (1.10) to both log-concave and $s$-concave functions, which are stated below as the following two theorems.
@@
Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f$
for $1<\alpha<\beta<\infty$. In each inclusion, there is equality if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$.
When $\alpha=n$, the inclusion
$\frac{1}{\Gamma(n+1)^{1/n}}{\rm R}_{n}f\subset 2\|f\|_{1}|\Pi^{*}f|,$
recovers (1.11). The inclusion relations in Theorem 4 can be deduced from the results of Alonso-Gutiérrez, Bernués and González Merino [2, Lemma 3.2, Remark 5], but the equality cases were not addressed there. In Section 9, we employ a different approach that allows us to establish Theorem 4 together with a complete characterization of equality.
In proving Theorem 4, we first establish an inclusion relation between ${\rm R}_{\alpha}f$ and ${\rm R}_{\beta}f$, and then consider the limits $\alpha\to-1$ and $\beta\to\infty$ to obtain inclusions between ${\rm R}_{\infty}f$ and $\Pi^{*}f$. However, in this limiting process $\beta\to\infty$ degenerates to the singular set $\{o\}$, so that ${\rm R}_{\infty}f$ does not appear in the general case. This motivates us to consider the embedding $K\mapsto\operatorname{1}_{K}$, leading to the following natural question:
Open problem. For a log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, does
hold? Here $c_{n,\alpha}$ is the constant in (1.10).
Unfortunately, this inclusion fails for general log-concave integrable functions with compact support, as we show by a counterexample in Section 9. This failure leads us naturally to consider $s$-concave functions.
Our main result concerns $s$-concave functions $f\in L^{1}(\mathbb{R}^{n})$, where we denote the positive part of $f$ by $f_{+}(x):=\max\{f(x),0\}$. The equality cases in this setting are fully characterized as well.
Theorem 5.
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. For $1<\alpha<\beta<\infty$,
where $c_{n,\alpha}(s)=(nB(\alpha+1,n+\frac{1}{s}))^{-1/\alpha}$. In each inclusion, there is equality if and only if $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$, where $c>0$, $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
Theorems 4 and 5 share the same general structure and rely on similar methods of proof. As noted above, in Theorem 4, $\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f$ degenerates to the singular set $\{o\}$ as $\beta\to\infty$, so that the set ${\rm R}_{\infty}f$ does not appear. In contrast, Theorem 5 avoids this degeneration, and in the limit $s\to\infty$, the inclusion (1.12) recovers (1.10).
2. Preliminary
We collect some basic notation and results on $L^{p}$ spaces, Schwarz symmetrals, star-shaped sets, log-concave functions, radial mean bodies and fractional polar projection bodies. The books of Gardner [14], Schneider [45], and Schneider and Weil [46] are good general references.
@@ 2.1. $L^{p}$ spaces
Let p>1p>-1 and $(X,\mathcal{B},\mu)$ be a measure space. In this paper, we mainly consider the case $X=\mathbb{R}^{n}$ or $X\subset\mathbb{R}^{n}$ with $\mu$ being the $n$-dimensional Lebesgue measure.
@@
$\|f\|_{p}=\Big(\int_{X}|f(x)|^{p}d\mu(x)\Big)^{1/p}.$
In particular, if $p\in(-1,0)$ and $\mu(\{x\in X:f(x)=0\})>0$, we set $\|f\|_{p}=+\infty$. The space $L^{p}(X)$ is then defined as
@@
where the $f=g$ in $L^{p}(X)$ means that they are equal almost everywhere.
When $\mu(x)=1$, the $L^{p}$-norm is monotone in $p$, that is,
@@
for 1<p<q<0-1<p<q<0 or $0<p<q<\infty$. This follows from Jensen’s inequality. In fact, the monotonicity also holds for $p<0<q$. To see this, under the assumption $\mu(X)=1$, we first note the limit
@@ 2.2. Symmetrization
For $E\subset\mathbb{R}^{n}$, the characteristic function $\operatorname{1}_{E}$ is defined by $\operatorname{1}_{E}(x)=1$ for $x\in E$, and $\operatorname{1}_{E}(x)=0$ for $x\notin E$. If $E$ is a Borel set of finite measure, the Schwarz symmetral of $E$, denoted by $E^{\star}$, is the closed, centered ball with the same volume as $E$.
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a non-negative measurable function. The superlevel set of $f$ is defined by $\{f\geq t\}=\{x\in X:f(x)\geq t\}$ for $t>0$. We say $f$ is non-zero if $\{f\neq 0\}$ has positive measure. For $f$ with superlevel sets of finite measure, by the layer cake formula,
@@
for almost every $x\in\mathbb{R}^{n}$.
The Schwarz symmetral of $f$, denoted by $f^{\star}$, is defined as
@@
for $x\in\mathbb{R}^{n}$. Hence $f^{\star}$ is radially symmetric and the superlevel set $\{f^{\star}\geq t\}$ has the same volume as $\{f\geq t\}$ for every $t>0$.
One main technique in our proofs is the Riesz rearrangement inequality (see, for example [31, Theorem 3.7]) and the characterization of equality case due to Burchard [8].
Theorem 2.1 (Riesz’s rearrangement inequality.).
@@ 2.3. Star-shaped sets and dual mixed volumes
We say $K\subset\mathbb{R}^{n}$ is a star-shaped set (with respect to the origin) if for any $x\in K$ the line segment $[0,x]$ is contained in $K$. The gauge function $\|\cdot\|_{K}:\mathbb{R}^{n}\rightarrow[0,+\infty]$ of a star-shaped set $K$ is defined by
$\|x\|_{K}=\inf\{\lambda>0:x\in\lambda K\},$
@@
$\rho_{K}(x)=\|x\|_{K}^{-1}=\sup\{\lambda\geq 0:\lambda x\in K\}.$
In particular, when $K$ is the unit ball, $\|\cdot\|_{K}$ is the Euclidean norm $|\cdot|$.
For a star-shaped set $K$ with measurable radial function, the $n$-dimensional volume, or the $n$-dimensional Lebesgue measure of $K$ is given by
@@
where $du$ is the spherical Lebesgue measure. The dual mixed volume for star-shaped sets $K$ and $L$ by Lutwak [34] is given by
@@
For $\alpha\in(0,n)$, and star-shaped sets $K$ and $L$ of finite volume, the dual mixed volume inequality states that
@@
where equality holds if and only if $K$ and $L$ are dilates, that is, $\rho_{K}(u)=c\rho_{L}(u)$ for some $c\geq 0$ and almost all $u\in\mathbb{S}^{n-1}$. For $\alpha<0$ and $\alpha>n$, the dual mixed volume inequality states that
@@ 2.4. Log-concave functions
[2, Lemma 2.1] Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a log-concave and integrable function. Then the function $g:\mathbb{R}^{n}\rightarrow[0,\infty)$ defined by
@@
The following theorem from Balogh and Kristály [6] gives the equality characterization for Borell-Brascamp-Lieb inequality.
@@
[6, Theorem 3.1] Let $\lambda\in(0,1)$ and $s>-\frac{1}{n}$. Let $f,g,h:\mathbb{R}^{n}\rightarrow[0,\infty)$ be three non-zero, compactly supported integrable functions satisfying
@@
$\int_{\mathbb{R}^{n}}h\geq{\mathcal{M}}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}f,\int_{\mathbb{R}^{n}}g\Big),$
@@
$\left\{\begin{aligned} &f(x)=\Phi(x)&&,\\ &g(c_{0}x+x_{0})=c_{0}^{\frac{1}{s}}\Phi(x)&&,\\ &h((1-\lambda+\lambda c_{0})x+\lambda x_{0})=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{0}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&.\end{aligned}\right$. (2.3)
2.5. Convex bodies and radial mean bodies
A convex body in $\mathbb{R}^{n}$ is a convex, compact set with non-empty interior. Let $K\subset\mathbb{R}^{n}$ be a convex body. For $\alpha>-1$ and $\alpha\neq 0$, Gardner and Zhang [18] defined the radial $\alpha$-th mean body of $K$ via its radial function by
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx.$
For $\alpha=0$,
$\log\rho_{{\rm R}_{0}f}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx.$
For $x=x^{\prime}+tu$ with $x^{\prime}\in K|u^{\perp}$, we have
$\int_{K}\rho_{K-x}(u)^{\alpha}dx=\int_{K|u^{\perp}}\int_{t_{1}}^{t_{2}}(t_{2}-t)^{\alpha}dtdx^{\prime},$
where $t_{1}=\min\{t:x^{\prime}+tu\in K\}$ and $t_{2}=\max\{t:x^{\prime}+tu\in K\}$. Since $t_{2}-t_{1}=|K\cap(x^{\prime}+\mathbb{R}u)|$, we then obtain
$\int_{K}\rho_{K-x}(u)^{\alpha}du=\frac{1}{\alpha+1}\int_{K|u^{\perp}}|K\cap(x^{\prime}+\mathbb{R}u)|^{\alpha+1}dx^{\prime}.$
@@ 2.6. Fractional polar projection bodies
For $\alpha\in(-1,0)$ and measurable $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$,
$\rho_{\Pi^{*,-\alpha}f}(u)^{-\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x+tu)-f(x)|dxdt.$ (2.4)
Then the affine fractional Sobolev inequality (1.7) can be written as
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|\Pi^{*,-\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy.$
Following the calculation in [27], we obtain that
$\rho_{\Pi^{*,-\alpha}\operatorname{1}_{K}}(u)^{\alpha}$ $=\int_{0}^{\infty}t^{\alpha-1}|K\triangle(K+tu)|dt$ (2.5)
$=\int_{0}^{\infty}t^{\alpha-1}\int_{K|u^{\perp}}2\min\{|K\cap(y+\mathbb{R}u),t|\}dydt$
$=2\int_{K|u^{\perp}}\int_{0}^{\infty}t^{\alpha-1}\min\{|K\cap(y+\mathbb{R}u)|,t\}dtdy$
$=2\int_{K|u^{\perp}}\Big(\int_{0}^{|K\cap(y+\mathbb{R}u)|}t^{\alpha}dt+\int_{|K\cap(y+\mathbb{R}u)|}^{\infty}|K\cap(y+\mathbb{R}u)|t^{\alpha-1}dt\Big)dy$
$=-\frac{2}{\alpha(\alpha+1)}\int_{K|u^{\perp}}|K\cap(y+\mathbb{R}u)|^{\alpha+1}dy,$
where $K\triangle L$ is the symmetric difference of $K,L\subset\mathbb{R}^{n}$. Hence,
$\Pi^{*,-\alpha}\operatorname{1}_{K}=\Big(\frac{2|K|}{-\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K,$ (2.6)
for $\alpha\in(-1,0)$.
3. Proof of Theorem 1
In this section, we focus on the proof of Theorem 1. Before presenting the proof, we first establish the following lemma, which is based on the Minkowski inequality (see, for example, [31, Theorem 2.4]). Let $F$ be a non-negative, measurable function on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ and $p\geq 1$. The Minkowski inequality states that
$\int_{\mathbb{R}^{n}}\Big(\int_{\mathbb{R}^{n}}F(x,y)^{p}dx\Big)^{1/p}dy\geq\Big(\int_{\mathbb{R}^{n}}\Big(\int_{\mathbb{R}^{n}}F(x,y)dy\Big)^{p}dx\Big)^{1/p}.$ (3.1)
Equality and finiteness in (3.1) for $p\in(1,\infty)$ imply the existence of measurable functions $\varphi,\psi:\mathbb{R}^{n}\rightarrow[0,\infty)$ such that $F(x,y)=\varphi(x)\psi(y)$ almost everywhere.
Lemma 3.1.
For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$\|f\|_{\frac{n}{n+\alpha}}\geq\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
Equality holds if and only if $f=c\operatorname{1}_{E}$ for some measurable $E$ of finite measure and some constant $c>0$.
Proof.
By the Minkowski inequality (3.1),
$\int_{\mathbb{R}^{n}}f(x)^{\frac{n}{n+\alpha}}dx$ $=\int_{\mathbb{R}^{n}}\Big(\int_{0}^{\infty}\operatorname{1}_{\{f\geq r\}}(x)^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}}dx$ (3.2)
$\geq\Big(\int_{0}^{\infty}\Big(\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)dx\Big)^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}}$
$=\Big(\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}},$
that is
$\|f\|_{\frac{n}{n+\alpha}}\geq\int_{0}^{\infty}|\{f>t\}|^{\frac{n+\alpha}{n}}dt.$
It is clear that equality is attained when $f=c\operatorname{1}_{E}$. Conversely, if equality holds in (3.2), there are non-negative measurable functions $\varphi$ and $\psi$ such that
$\operatorname{1}_{\{f\geq r\}}(x)=\varphi(x)\psi(t).$
This implies that
$\{(x,r)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)\geq r\}=\{x\in\mathbb{R}^{n}:\varphi(x)>0\}\times\{r\in\mathbb{R}:\psi(r)>0\}.$
Therefore, for every fixed $r\in\mathbb{R}$ such that $r>0$ and $r\leq f(x)$ for some $x$, the superlevel set $\{f\geq r\}$ is a set independent of $r$, which implies that $f=c\operatorname{1}_{E}$ for some measurable set $E$ of finite measure and some constant $c>0$. ∎
We are now ready to prove Theorem 1.
Proof of Theorem 1.
Note that
$\min\{f(x),f(y)\}=\int_{0}^{\infty}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)dr.$
By Fubini’s theorem and Riesz rearrangement inequality, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)}{|x-y|^{n-\alpha}}dxdydr$
$\leq\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{\{f\geq r\}^{\star}}(y)}{|x-y|^{n-\alpha}}dxdydr$
$=\sigma_{n,\alpha}\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr,$
where the last equation is from the definition of chord power integrals and $|E^{\star}|=|E|$ for any $E\subset\mathbb{R}^{n}$ with finite measure. The best constant $\sigma_{n,\alpha}$ is given by
$\sigma_{n,\alpha}=\frac{nI_{\alpha+1}(B^{n})}{\alpha(\alpha+1)\omega_{n}^{\frac{\alpha}{n}}}.$
Then Lemma 3.1 implies that
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
When the equality holds, it follows from Lemma 3.1 that $f$ is a characteristic function. Then by the isoperimetric inequality of chord integrals, we have $f=c\operatorname{1}_{B}$, up to a set of measure zero. ∎

### v2

@@ 1. Introduction
Using the second correspondence, Langharst, Marín Sola and Ulivelli [28] showed that for log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f.$ (1.13)
Equality holds in each inclusion if and only if $f(x)=\|f\|_{\infty}e^{-\|x-x_{0}\|_{\triangle}}$, where $x_{0}$ is a maximizer of $f$ and $\triangle$ is a simplex containing the origin and $x_{0}$. When $\alpha=n$, the right-most inclusion recovers Zhang’s inequality for log-concave functions, as established by Alonso-Gutiérrez, Bernués and González Merino [3]. However, in the limit $\beta\rightarrow\infty$, the body ${\rm R}_{\infty}f$ no longer appears in (1.13). This indicates a limitation of the second correspondence in capturing the endpoint behavior of radial mean bodies. On the other hand, the monotonicity property (1.11) naturally extends from convex bodies to functions via the embedding $K\mapsto\operatorname{1}_{K}$. These considerations motivate us to adopt the first correspondence and lead to the following question.
Question. For a non-zero, log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, does
hold for $1<\alpha<\beta<\infty$? Here $c_{n,\alpha}$ is the constant in (1.12).
Unfortunately, this inclusion fails for general log-concave functions, as we show by a counterexample in Section 6. This failure leads us to consider $s$-concave functions with $s>0$.
Let $f_{+}(x):=\max\{f(x),0\}$ denote its positive part. The following theorem extends (1.12) to $s$-concave functions, with fully characterized equality cases.
@@
Let $s>0$ and let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. For $1<\alpha<\beta<\infty$, the following chain of inclusions holds:
${\rm R}_{\infty}f\subset c_{n,\beta}(s)\,{\rm R}_{\beta}f\subset c_{n,\alpha}(s)\,{\rm R}_{\alpha}f\subset\Big(n+\frac{1}{s}\Big)\|f\|_{1}\,\Pi^{*}f,$ (1.14)
where $c_{n,\alpha}(s)=\bigl((n+\tfrac{1}{s})\,B(\alpha+1,n+\tfrac{1}{s})\bigr)^{-1/\alpha}$. In each inclusion, there is equality if and only if $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
Letting $\beta=n$ and $\alpha=n$ in (1.14) yields the Rogers-Shephard inequality and Zhang inequality, respectively, for $s$-concave functions; see Section 6 for detailed statement.
The optimizer in (1.14) admits a simple geometric interpretation. Let $\triangle$ be a simplex in $\mathbb{R}^{n}$ with vertices $v_{0},\ldots,v_{n}$, and identify $\mathbb{R}^{n}$ with the hyperplane $\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n+1}$. For $f^{s}(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}\quad\text{with }x_{0}\in\triangle$, its hypograph is given by the set
$\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm supp}(f^{s}),~0\leq l\leq f^{s}(x)\}$
and coincides exactly with the simplex in $\mathbb{R}^{n+1}$ having vertices $v_{0},\ldots,v_{n},(x_{0},a)$. The equality characterization of (1.14) relies on the equality case of the Borell–Brascamp–Lieb inequality. Its highly rigid structure accounts for the difficulty in the study of the equality cases.
2. Preliminaries
In this section, we collect basic notation and results on function spaces and Schwarz symmetrals, together with some basic geometric facts. Useful general references include the books by Gardner [16], Schneider [42], and Schneider and Weil [43].
@@ 2.1. $L^{p}$ spaces
Let p>1p>-1 and $(X,\mathcal{B},\mu)$ be a measure space. In this work, we primarily focus on the cases $X=\mathbb{R}^{n}$ or $X\subset\mathbb{R}^{n}$ with $\mu$ being the $n$-dimensional Lebesgue measure.
@@
$\|f\|_{p}=\Big(\int_{X}|f(x)|^{p}d\mu(x)\Big)^{1/p},$
and
$\|f\|_{\infty}=\inf\{C>0:|f(x)|\leq C~~\text{almost everywhere on}~X\}.$
In particular, for $p\in(-1,0)$, we set $\|f\|_{p}=\infty$ if $\mu(\{x\in X:f(x)=0\})>0$. The associated space $L^{p}(X)$ is then defined as
@@
where $f=g$ in $L^{p}(X)$ means that they are equal almost everywhere.
When $\mu(X)=1$, the $L^{p}$ norm is monotone in $p$: for 1<p<q<0-1<p<q<0 or $0<p<q<\infty$,
@@
as a consequence of Jensen’s inequality. In fact, this monotonicity also extends for $p<0<q$. To see this, under the assumption $\mu(X)=1$, we first note the limit
@@ 2.2. Symmetrization
For a Borel set $E\subset\mathbb{R}^{n}$ with finite measure, the Schwarz symmetral of $E$, denoted by $E^{\star}$, is the closed, centered ball having the same volume as $E$. Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. For $t>0$, the superlevel set of $f$ is defined by $\{f\geq t\}=\{x\in\mathbb{R}^{n}:f(x)\geq t\}$. We say $f$ is non-zero if $\{f\neq 0\}$ has positive measure.
If the superlevel sets of $f$ have finite measure, the layer cake formula gives
@@
for almost every $x\in\mathbb{R}^{n}$. The Schwarz symmetral of $f$, denoted by $f^{\star}$, is defined as
@@
for $x\in\mathbb{R}^{n}$, so that $f^{\star}$ is radially symmetric and the superlevel sets satisfy
$|\{f^{\star}\geq t\}|=|\{f\geq t\}|$
for all $t>0$.
A central tool in our proofs is the Riesz rearrangement inequality (see, for example [29, Theorem 3.7]) together with the characterization of equality by Burchard [9].
Theorem 2.1 (Riesz’s rearrangement inequality).
@@ 2.3. Star-shaped sets and dual mixed volumes
A set $K\subset\mathbb{R}^{n}$ is called star-shaped (with respect to the origin) if for every $x\in K$, the line segment $[o,x]$ is contained in $K$. For a star-shaped set $K$ containing the origin in its interior, the gauge function $\|\cdot\|_{K}:\mathbb{R}^{n}\rightarrow[0,\infty]$ is defined by
$\|x\|_{K}=\inf\{\lambda>0:x\in\lambda K\}.$ (2.1)
If $K$ contains the origin, possibly on its boundary, then for any point $x$, the gauge function $\|x\|_{K}$ is defined by (2.1) if there exists $r>0$ such that $rx\in K$; otherwise, $\|x\|_{K}=\infty$.
@@
$\rho_{K}(x)=\|x\|_{K}^{-1}=\sup\{r\geq 0:rx\in K\}.$
In particular, if $K$ is the unit Euclidean ball, $\|\cdot\|_{K}$ coincides with the Euclidean norm $|\cdot|$.
For a star-shaped set $K$ with measurable radial function, the volume, or the $n$-dimensional Lebesgue measure of $K$ is given by
@@
where $du$ denotes the spherical Lebesgue measure. Following Lutwak [32], the dual mixed volume for star-shaped sets $K$ and $L$ is defined by
@@
$\tilde{V}_{\log}(K,L)=\lim_{\alpha\rightarrow 0}\log\Big(\frac{\tilde{V}_{\alpha}(K,L)}{|K|}\Big)^{1/\alpha}=\frac{1}{n|K|}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n}\log\Big(\frac{\rho_{K}(u)}{\rho_{L}(u)}\Big)du.$
If $\alpha\in(0,n)$ and $K,L$ are star-shaped sets with finite volume, the dual mixed volume inequality states that
@@
with equality holds if and only if $K$ and $L$ are dilates, i.e., $\rho_{K}(u)=c\rho_{L}(u)$ for some $c\geq 0$ and almost all $u\in\mathbb{S}^{n-1}$. For $\alpha<0$ or $\alpha>n$, the dual mixed volume inequality states that
@@ 2.4. Log-concave and $s$-concave functions
[3, Lemma 2.1] Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a log-concave and integrable function. Then the function $g:\mathbb{R}^{n}\rightarrow[0,\infty)$ defined by
@@
The following theorem of Balogh and Kristály [6] gives the equality characterization of the Borell-Brascamp-Lieb inequality.
@@
[6, Theorem 3.1] Let $\lambda\in(0,1)$ and $s>-\frac{1}{n}$. Let $F,G,H:\mathbb{R}^{n}\rightarrow[0,\infty)$ be non-zero, compactly supported integrable functions satisfying
@@
$\left\{\begin{aligned} &F(x)=\Phi(x)&&,\\ &G(c_{0}x+x_{0})=c_{0}^{\frac{1}{s}}\Phi(x)&&,\\ &H((1-\lambda+\lambda c_{0})x+\lambda x_{0})=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{0}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&.\end{aligned}\right$.
@@ 2.5. Fractional polar projection bodies
For $\alpha\in(-1,0)$ and measurable $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, the fractional polar projection body $\Pi^{*,-\alpha}f$ is defined by
$\rho_{\Pi^{*,-\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x)-f(x+ru)|dxdr.$
The affine fractional Sobolev inequalities (1.1) state that
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|\Pi^{*,-\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy$
for $f\in W^{-\alpha,1}(\mathbb{R}^{n})$
Let $\mathcal{B}(\mathbb{R}^{n})$ denotes the class of Borel set in $\mathbb{R}^{n}$. Let $C_{c}^{\infty}(\mathbb{R}^{n};\mathbb{R}^{n})$ denote the set of smooth vector fields with compact support. Moreover, ${\rm div}\,\nu$ denotes the divergence of $\nu\in C_{c}^{\infty}(\mathbb{R}^{n};\mathbb{R}^{n})$. We say that $f\in L^{1}(\mathbb{R}^{n})$ is a function of bounded variation on $\mathbb{R}^{n}$ if there is a finite vector-valued Radon measure $\mathcal{D}f:\mathcal{B}(\mathbb{R}^{n})\rightarrow\mathbb{R}^{n}$ such that
$\int_{\mathbb{R}^{n}}\nu(x)\cdot n_{f}(x)d|\mathcal{D}f|(x)=-\int_{\mathbb{R}^{n}}f(x){\rm div}\,\nu(x)dx$
for every $\nu\in C_{c}^{\infty}(\mathbb{R}^{n};\mathbb{R}^{n})$, where $|\mathcal{D}f|:\mathcal{B}(\mathbb{R}^{n})\rightarrow[0,\infty)$ denotes the variation measure of $\mathcal{D}f$ and $n_{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ the Radon-Nikodym derivative of $\mathcal{D}f$ with respect to $|\mathcal{D}f|$.
For $f\in L^{1}(\mathbb{R}^{n})$ with bounded variation, the polar projection body $\Pi^{*}f$ is defined by
$\|u\|_{\Pi^{*}f}=\frac{1}{2}\int_{\mathbb{R}^{n}}|n_{f}(x)\cdot u|\,d|\mathcal{D}f|(x).$
For more information about $\Pi^{*}f$, we refer the reader to [44].